2 cars leave at a town at the same time & travel in opposite directions .the speed of one car is 12 km/hr moere than the other . thjey are 444 km apart after 3 hours . find the speed of each car .
Answer by checkley77(12844)

D=RT
D1=R3 FOR CAR ONE.
D2=(R+12)3 FOR CAR TWO.
D1+D2=444
444=R*3
R=444/3
R=148
R+(R+12)=148
2R=148-12
2R=136
R=136/2
R=68 KM/HR. IS THE SPEED OF THE SLOWER CAR.
68+12=80 KM/HR IS THE SPEED OF THE FASTER CAR.
PROOF;
444=68*3+80*3
444=204+240
444=444